# Efficient Reinforcement Finetuning via Adaptive Curriculum Learning

### v1

@@ 1 Introduction
Reinforcement Finetuning (RFT) has emerged as a powerful technique for aligning large language models (LLMs) with task-specific goals, particularly in domains such as mathematics and code generation where correctness is well defined (DeepSeek-AI et al., 2025; OpenAI et al., 2024). By optimizing a policy model with reward signals that reflect task success, RFT enables more targeted learning than supervised finetuning (SFT) alone. However, despite its promise, RFT remains sample-inefficient and computationally expensive. Its training involves repeated rollout generation, reward computation, and policy updates—making it costly and difficult to scale (Ahmadian et al., 2024; Kazemnejad et al., 2024; Li et al., 2024; Hu, 2025; Cui et al., 2025). Recent efforts to address RFT inefficiency have focused on algorithmic simplification (e.g., RAFT (Dong et al., 2023), GRPO (DeepSeek-AI et al., 2025), ReMax (Li et al., 2024)), and data-centric strategies (e.g., LIMO (Ye et al., 2025), LIMR (Li et al., 2025)). While these approaches improve sample or compute efficiency, they often introduce trade-offs: algorithmic simplifications may increase variance or limit stability, and static data filtering or scoring can be brittle, computationally heavy, or model-specific. Moreover, most methods’ success relies on fixed datasets or training schedules, which can be suboptimal in non-uniform or imbalanced data regimes.
To overcome these limitations, we propose AdaRFT, a reinforcement finetuning method based on adaptive curriculum learning (Bengio et al., 2009), which dynamically adjusts training set difficulty to match the model’s evolving skill level. The algorithm is shown in Alg. 1. To illustrate the motivation, consider asking a grade school student to solve a problem from the International Mathematical Olympiad (IMO). Without the necessary background and tools, the student would struggle to engage meaningfully with the task, let alone learn from it. Inspired by this analogy, we hypothesize that LLMs trained via RFT may face a similar challenge: when presented with tasks that exceed their current capabilities, the learning signal becomes weak or ineffective. On the other hand, tasks that are too simple may not sufficiently challenge the model to improve. In both scenarios, a mismatch between task difficulty and model ability leads to inefficient learning.
AdaRFT addresses this challenge by dynamically selecting training problems that match the model’s current skill level. Each problem in the dataset is preassigned a difficulty score, which can be either estimated automatically or derived from human annotations. During training, we maintain a target difficulty level that is adjusted based on the model’s recent reward signal: as the model improves, the target difficulty increases; if performance drops, it decreases. At each step, the model is finetuned on problems whose difficulty is closest to this target, ensuring a steady progression through increasingly challenging but solvable tasks.
@@
Unlike prior work that simplifies the RL algorithm or precomputes model-specific learning-impact metrics, AdaRFT adds only a lightweight curriculum scheduler on top of standard RFT algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017), with no modification to reward functions and no model-dependent preprocessing. While AdaRFT is compatible with a variety of policy optimization methods (e.g., PPO, GRPO), in this work, we instantiate it using PPO for all experiments to demonstrate its effectiveness. We denote this specific instantiation as AdaRFT (PPO). We evaluate AdaRFT on a dataset spanning a wide range of competition-level math problems, including AMC, AIME, and IMO-style questions. Across multiple training distributions and two model sizes, adaptive curriculum learning significantly improves both training efficiency and final performance. Gains are especially notable in imbalanced data regimes, where static sampling often fails. AdaRFT can speed up training by up to 2, offering a practical and scalable path to more efficient RFT in structured reasoning tasks.
@@ 2 Related Work
Most RFT pipelines are based on Proximal Policy Optimization (PPO) (Schulman et al., 2017), with recent variants such as RAFT (Dong et al., 2023), GRPO (DeepSeek-AI et al., 2025), REINFORCE++ (Hu, 2025), and ReMax (Li et al., 2024) aiming to reduce computational cost by simplifying the underlying RL algorithm. These methods generally improve efficiency and stability, but often introduce trade-offs. For example, GRPO eliminates the value function by estimating advantages through relative comparisons within groups of outputs. While this removes the need for a critic and reduces complexity, it can lead to instability due to noisier advantage estimates, higher variance in updates, and increased sample requirements (DeepSeek-AI et al., 2025; Hu, 2025). In parallel, data-centric strategies have emerged as promising alternatives for efficient finetuning. Techniques like LIMO (Ye et al., 2025) and s1 (Muennighoff et al., 2025) demonstrate that small, carefully curated supervised finetuning datasets can achieve strong performance while using orders of magnitude less data. However, they rely heavily on effective data selection and prompt design, which may hinder generalization to new tasks. LIMR (Li et al., 2025) also aims to reduce RFT sample requirements by scoring training examples using a learning impact measurement (LIM). Although this allows subsequent training with fewer samples, computing LIM scores requires a full training run on the entire dataset. Moreover, despite using fewer samples afterwards, models still take roughly the same number of steps to converge, limiting the overall efficiency gains. LIMR is also model-dependent, meaning that LIM scores must be recomputed when switching to a different model, further increasing overhead. In contrast to these approaches, AdaRFT introduces an adaptive curriculum learning strategy that adaptively selects training problems based on the model’s current performance. Rather than simplifying the RL algorithm or relying on static data selection, we continuously adjust the training difficulty to match the model’s evolving capability. This curriculum-aware training leads to faster convergence, improved accuracy, and higher sample and compute efficiency, especially in non-uniform training distributions, while maintaining stability and generalization across tasks.
@@
Curriculum learning (CL) (Bengio et al., 2009) is a training strategy that structures learning by presenting training samples in a meaningful progression, typically from simpler to more difficult tasks. In RL, task-specific curricula sort tasks based on predefined or learned difficulty metrics to guide agents through progressively harder challenges (Zaremba & Sutskever, 2015; Justesen et al., 2018; Wang et al., 2019). Teacher-guided curricula treat task selection as a bandit or partially observable Markov decision process problem (POMDP), where a teacher agent adaptively assigns tasks that maximize the student’s learning progress (Matiisen et al., 2017; Portelas et al., 2019). Self-play methods generate an automatic curriculum by having agents propose and solve tasks for each other, dynamically increasing task difficulty through competition (Sukhbaatar et al., 2018). Goal generation techniques, such as Goal GANs, produce goals of intermediate difficulty based on current policy performance to optimize learning in sparse-reward settings (Florensa et al., 2018). Skill-based curricula organize unsupervised trajectories into a latent skill space to create transferable training tasks for meta-RL (Jabri et al., 2019). Distillation-based approaches, like progressive networks (Rusu et al., 2022) and Mix & Match (Czarnecki et al., 2018), transfer knowledge from simpler agents to more complex ones through shared architectures and imitation losses. These diverse approaches demonstrate the value of structured task progression in improving generalization and sample efficiency across complex RL domains. While curriculum learning has been widely studied in classical RL, its application to RFT of LLMs remains largely unexplored. Our work is among the first to introduce adaptive curriculum learning in the RFT setting, adapting problem difficulty in real-time based on model performance. This approach improves both sample efficiency and final performance, demonstrating that curriculum learning is a powerful tool for scaling RFT on complex reasoning tasks.
@@ 3.1 Dynamic Curriculum Sampling
To construct an adaptive curriculum, we define a target difficulty , which represents the current target difficulty level for training. AdaRFT dynamically adjusts based on the model’s reward signal to maintain an optimal difficulty level for learning.
At each training step (see Algorithm 1), the algorithm computes the absolute difference between the target difficulty and the difficulty of each problem in the dataset (Step 1.4): for all . The batch of training problems is formed by selecting the problems with the smallest values of (Step 1.5), producing a batch: .
This ensures that the selected problems are closest to the model’s current target difficulty, focusing the learning process on problems that are neither too easy nor too hard.
@@ 3.2 Policy Update
The selected batch is used to train the policy model , which generates responses (Step 1.6): . A reward signal is then computed based on the correctness of the model’s output (Step 1.7):
The average reward over the batch is computed as (Step 1.7): .
The policy can then be updated using a reinforcement learning algorithm such as PPO, GRPO, or REINFORCE++ (Step 1.8): .
@@ 3.3 Target Difficulty Update
To adapt the curriculum dynamically, the target difficulty is updated based on the average reward. If the model performs well on the current difficulty level (high reward), the target difficulty increases, making the training problems harder. Conversely, if the model performs poorly, the target difficulty decreases. This dynamic update mechanism lies at the core of AdaRFT’s curriculum adaptation strategy.
The update rule (Step 1.9) is defined as:
Here, are hyperparameters: is the step size for adjusting the target difficulty, controls the sensitivity of the update, and is the target reward level, representing the desired success rate. The function ensures smooth updates and prevents large jumps in difficulty by saturating for large deviations, while the “clip” function constrains the target difficulty within the valid range . These bounds can be manually specified or automatically derived from the training set, for example by taking the minimum and maximum of the difficulty scores . Intuition and guidance for selecting these hyperparameters are discussed in Section 4.3.
@@ 4.1 Difficulty Estimation
Accurate estimation of problem difficulty is critical for AdaRFT. For difficulty estimation, we use the Qwen 2.5 MATH 7B model (Qwen et al., 2025).111https://huggingface.co/Qwen/Qwen2.5-Math-7B A model that is too strong (e.g., OpenAI o1 (OpenAI et al., 2024), DeepSeek R1 (DeepSeek-AI et al., 2025)) would solve most problems on the first attempt, leading to poor discrimination between easy and hard problems. Conversely, a model that is too weak (e.g., LLaMA 3.3 1B (Grattafiori et al., 2024)) would fail to solve most problems even after multiple attempts, limiting the signal required for curriculum adaptation.
We select Qwen 2.5 MATH 7B as the estimator because it demonstrates a balanced solving ability — it can solve moderately difficult problems while still failing on more complex problems, thereby providing a more informative difficulty signal. For each problem, the difficulty score is computed as:
where is the number of attempts allowed for the model. In our setup, we use . This approach ensures that difficulty scores reflect the true problem-solving capability of a moderately capable model, preventing over- or under-estimation of difficulty.
@@
To further validate the reliability of our difficulty estimates, we examined their alignment with the human-labeled difficulty levels provided in the MATH dataset. The MATH dataset comprises 12,500 competition-level mathematics problems sourced from contests such as the AMC 10, AMC 12, and AIME. Each problem is categorized into one of five difficulty levels, following the classification system used by the Art of Problem Solving (AoPS) community.222https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki%3ACompetition_ratings In this dataset, a subject’s easiest problems are assigned a difficulty level of 1, while the hardest problems are assigned a difficulty level of 5. For instance, the initial problems in an AMC 8 exam are often labeled as level 1, whereas AIME problems are typically labeled as level 5. As shown in Figure 1(b), there is a clear downward trend in average solved percentage as the labeled difficulty level increases, ranging from 86.0% at level 1 to 52.7% at level 5. This strong negative correlation indicates that the model’s empirical performance is consistent with the intended difficulty stratification, reinforcing the usefulness of both the labeled difficulty levels and our estimation approach for guiding curriculum learning.
@@ 4.2 Dataset
We use the DeepScaleR dataset (Luo et al., 2025) as the training set. DeepScaleR compiles problems from multiple sources, including the American Invitational Mathematics Examination (AIME) from 1984 to 2023 and the American Mathematics Competitions (AMC) prior to 2023. In addition, the dataset includes problems from the Omni-MATH (Gao et al., 2024) and Still datasets (Team, 2025), which feature problems from various national and international math competitions. This results in a diverse and challenging training set, covering a wide range of mathematical domains and difficulty levels.
@@ 4.3 Training Setup
We trained two models on the three difficulty-based distributions of the DeepScaleR dataset described in Section 4.2: Qwen 2.5 7B333https://huggingface.co/Qwen/Qwen2.5-7B and Qwen 2.5 MATH 1.5B444https://huggingface.co/Qwen/Qwen2.5-Math-1.5B. This setup allows us to evaluate the effectiveness of AdaRFT on models with different initial performance levels when exposed to skew-hard, skew-easy, and uniform problem distributions. Both models were trained using two approaches: the standard PPO algorithm and AdaRFT (PPO), our method that integrates adaptive curriculum learning with PPO (see Section 3). For AdaRFT, the target difficulty was dynamically adjusted throughout training based on the model’s reward signal. This adjustment ensured that the model consistently encountered problems that matched its current skill level, preventing the model from being overwhelmed by overly difficult problems or stagnating on problems that were too easy.

### v2

@@ 1 Introduction
Reinforcement Finetuning (RFT) has emerged as a powerful technique for aligning large language models (LLMs) with task-specific goals, particularly in domains such as mathematics and code generation where correctness is well defined (DeepSeek-AI et al., 2025; OpenAI et al., 2024b). By optimizing a policy model with reward signals that reflect task success, RFT enables more targeted learning than supervised finetuning (SFT) alone. However, despite its promise, RFT remains sample-inefficient and computationally expensive. Its training involves repeated rollout generation, reward computation, and policy updates—making it costly and difficult to scale (Ahmadian et al., 2024; Kazemnejad et al., 2024; Li et al., 2024; Hu, 2025; Cui et al., 2025). Recent efforts to address RFT inefficiency have focused on algorithmic simplification (e.g., RAFT (Dong et al., 2023), GRPO (DeepSeek-AI et al., 2025), ReMax (Li et al., 2024)), and data-centric strategies (e.g., LIMO (Ye et al., 2025), LIMR (Li et al., 2025)). While these approaches improve sample or compute efficiency, they often introduce trade-offs: algorithmic simplifications may increase variance or limit stability, and static data filtering or scoring can be brittle, computationally heavy, or model-specific. Moreover, most methods’ success relies on fixed datasets or training schedules, which can be suboptimal in non-uniform or imbalanced data regimes. More recently, early efforts have introduced curriculum-like ideas into RFT. Staged curricula divide training into a few manually-defined phases of increasing difficulty (Wen et al., 2025; Luo et al., 2025; Song et al., 2025), but these are coarse-grained and lack adaptivity. Other methods use online data filtering, repeatedly rolling out and pruning training samples until the model’s average reward meets a target threshold (Bae et al., 2025; Yu et al., 2025). While this approach helps prevent the model from stagnating on problems that are either too easy or too difficult, it is not truly adaptive and incurs significant rollout overhead.
@@
To address these limitations, we propose AdaRFT, a reinforcement finetuning method based on adaptive curriculum learning (Bengio et al., 2009), which dynamically adjusts training set difficulty to match the model’s evolving skill level. The intuition is simple: learning is most effective when tasks are neither too easy nor too hard. AdaRFT formalizes this by maintaining a target difficulty level, which increases or decreases based on recent reward feedback. At each step, the model is trained on examples closest to this target, promoting a steady progression through solvable yet challenging tasks. The full algorithm is outlined in Algorithm 1. Unlike prior work that relies on fixed stages, repeated rollouts, or model-specific data processing, AdaRFT is lightweight, general, and model-agnostic. It can be directly applied on top of any standard reinforcement learning (RL) algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017b). We evaluate AdaRFT on a dataset spanning a wide range of competition-level math problems, including AMC, AIME, and IMO-style questions. Across multiple training distributions and two model sizes, AdaRFT significantly improves both training efficiency and final performance. Gains are especially notable in imbalanced data regimes, where static sampling often fails. AdaRFT can reduce training time by up to 2, offering a practical and scalable path to more efficient RFT in structured reasoning tasks.
@@ 2 Related Work
Most RFT pipelines build on Proximal Policy Optimization (PPO) (Schulman et al., 2017b), with recent variants like RAFT (Dong et al., 2023), ReMax (Li et al., 2024), GRPO (DeepSeek-AI et al., 2025), and REINFORCE++ (Hu, 2025), aiming to reduce computational overhead by simplifying RL components. While effective, these methods often trade off stability or sample efficiency. In parallel, data-centric strategies have emerged as promising alternatives for efficient finetuning. LIMO (Ye et al., 2025) and s1 (Muennighoff et al., 2025) show that small, carefully selected supervised datasets can yield strong downstream performance, but their success hinges on manual curation, prompt engineering, and careful dataset construction, which may not generalize across tasks or models. LIMR (Li et al., 2025) and Wang et al. (2025) proposes scoring training examples based on their estimated learning impact, enabling selective finetuning with fewer samples. Yet, computing these scores requires a full training run, and the scores must be recomputed for each new model, limiting practicality and scalability. In contrast, AdaRFT introduces a lightweight, model-agnostic curriculum learning strategy that dynamically adjusts task difficulty based on reward feedback. This allows continuous adaptation to the model’s capabilities, improving convergence speed and final accuracy without modifying the RL algorithm or requiring manual data curation.
@@
Curriculum learning (CL) structures training by presenting tasks in an organized progression, typically from easy to hard, to enhance learning efficiency and generalization (Bengio et al., 2009). In RL, CL methods include task sorting by difficulty (Zaremba & Sutskever, 2015; Justesen et al., 2018; Wang et al., 2019), teacher-student frameworks that adaptively select tasks based on learning progress (Matiisen et al., 2017; Portelas et al., 2019), and self-play approaches that induce automatic curricula through agent competition (Sukhbaatar et al., 2018). Other strategies use intermediate-goal generation in sparse-reward settings (Florensa et al., 2018), unsupervised skill discovery (Jabri et al., 2019), or knowledge transfer via progressive networks and imitation (Czarnecki et al., 2018; Rusu et al., 2022). While CL is well-studied in classical RL, its application to RFT of LLMs is still limited. Existing methods typically use staged training with hand-designed difficulty tiers (Wen et al., 2025; Luo et al., 2025; Song et al., 2025), or online filtering schemes that repeatedly sample and discard data until rewards reach a target range (Bae et al., 2025; Yu et al., 2025). These methods either lack adaptability or introduce significant computational overhead due to repeated rollouts. In contrast, AdaRFT is among the first truly adaptive curriculum learning approaches for RFT: it continuously adjusts task difficulty based on the model’s reward signal, enabling efficient, scalable training without fixed schedules or repeated rollouts.
@@ 3.1 Dynamic Curriculum Sampling
To construct an adaptive curriculum, we define a target difficulty , which represents the current target difficulty level for training. AdaRFT dynamically adjusts based on the model’s reward signal to maintain an optimal difficulty level for learning. At each training step (see Alg. 1), the algorithm computes the absolute difference between the target difficulty and the difficulty of each problem in the dataset (Alg. 1, line 4): for all . The batch of training problems is formed by selecting the problems with the smallest values of (Alg. 1, line 5), producing a batch: . This ensures that the selected problems are closest to the model’s current target difficulty, focusing the learning process on problems that are neither too easy nor too hard.
@@ 3.2 Policy Update
The selected batch is used to train the policy model , which generates responses: . A reward signal is computed based on the correctness of the model’s output (Alg. 1, line 7):
@@ 3.3 Target Difficulty Update
To adapt the curriculum dynamically, the target difficulty is updated based on the average reward. If the model performs well on the current difficulty level (high reward), the target difficulty increases, making the training problems harder. Conversely, if the model performs poorly, the target difficulty decreases. This dynamic update mechanism lies at the core of AdaRFT’s curriculum adaptation strategy. The update rule (Alg. 1, line 9) is defined as:
Here, are hyperparameters: is the step size for adjusting the target difficulty, controls the sensitivity of the update, and is the target reward level, representing the desired success rate. The function ensures smooth updates and prevents large jumps in difficulty by saturating for large deviations, while the “clip” function constrains the target difficulty within the valid range . These bounds can be manually specified or automatically derived from the training set, for example by taking the minimum and maximum of the difficulty scores . Intuition and guidance for selecting these hyperparameters are discussed in Section 3.4 and 4.3.
3.4 Theoretical Justification for Target Reward
A key component of AdaRFT is its adaptive curriculum mechanism, which steers training toward a target reward level . Intuitively, we aim to train on examples that are neither trivially easy nor prohibitively hard. In this light, setting , corresponding to a success rate of roughly 50%, naturally aligns with this goal. This section formalizes that intuition by analyzing the relationship between reward variance and learnability in RFT with binary rewards.
In entropy-regularized reinforcement learning, the optimal policy can be expressed relative to a reference policy as (Korbak et al., 2022; Go et al., 2023; Rafailov et al., 2023):
where is the inverse temperature parameter controlling entropy regularization, and is the partition function that normalizes the action probability. The corresponding optimal value function and the partition function is given by (Schulman et al., 2017a; Richemond et al., 2024):
We can then take the expectation of the log-ratio between the optimal policy and the initial policy with respect to , leading to (Haarnoja et al., 2017; Schulman et al., 2017a):
Since the left-hand side can be interpreted as the negative reverse KL divergence between and (Rafailov et al., 2024), Bae et al. (2025) show that when the reward with is Bernoulli, the KL divergence is lower-bounded by the reward variance:
where is the model’s success rate on prompt . This implies that the lower bound on the KL divergence, and consequently the gradient magnitude during policy updates, is proportional to the reward variance, which is maximized when . In other words, training on prompts that the model succeeds on roughly half the time may yield the strongest learning signal.
@@ 4.1 Difficulty Estimation
Accurate estimation of problem difficulty is critical for AdaRFT. For difficulty estimation, we use the Qwen 2.5 MATH 7B model (Qwen et al., 2025).111https://huggingface.co/Qwen/Qwen2.5-Math-7B A model that is too strong (e.g., OpenAI o1 (OpenAI et al., 2024b), DeepSeek R1 (DeepSeek-AI et al., 2025)) would solve most problems on the first attempt, leading to poor discrimination between easy and hard problems. Conversely, a model that is too weak (e.g., LLaMA 3.3 1B (Grattafiori et al., 2024)) would fail to solve most problems even after multiple attempts, limiting the signal required for curriculum adaptation.
We select Qwen 2.5 MATH 7B as the estimator because it demonstrates a balanced solving ability. It can solve moderately difficult problems while still failing on more complex problems, thereby providing a more informative difficulty signal. For each problem, the difficulty score is computed as:
where is the number of attempts per problem. In our setup, we use . This approach ensures that difficulty scores reflect the true problem-solving capability of a moderately capable model, preventing over- or under-estimation of difficulty.
@@
To further validate the reliability of our difficulty estimates, we examined their alignment with the difficulty levels provided in the MATH dataset. The MATH dataset comprises 12,500 competition-level mathematics problems sourced from contests such as AMC 10, AMC 12, and AIME. Each problem is categorized into one of five difficulty levels, following the classification system used by the Art of Problem Solving (AoPS) community.222https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki%3ACompetition_ratings In this system, level 1 denotes the easiest problems, while level 5 represents the most difficult. For example, the initial problems in an AMC 8 exam are typically labeled as level 1, whereas AIME problems are often labeled as level 5. As shown in Figure 1(b), there is a clear downward trend in the average solve rate as the labeled difficulty level increases, ranging from 86.0% at level 1 to 52.7% at level 5. Specifically, the AoPS-derived difficulty levels yield a Pearson correlation of () with model success rates. This negative correlation indicates that the model’s empirical performance aligns well with the intended difficulty stratification, reinforcing the utility of both the labeled difficulty levels and our estimation approach in guiding curriculum learning. To further streamline the difficulty estimation process, we also prompted GPT-4o (gpt-4o-0806) (OpenAI et al., 2024a) to assign difficulty levels to the DeepScaleR dataset based on the AoPS rubric. Each problem was presented to GPT-4o with a request to rate its difficulty according to AoPS guidelines (the full prompt is shown in Appendix A.3). This approach provides a lightweight and scalable alternative to rollout-based estimation. As shown in Figure 1(b), GPT-4o’s difficulty ratings also correlate well with the model success rates, with a Pearson correlation of (), making it a practical proxy for curriculum scheduling when computational resources are constrained.
@@ 4.2 Dataset
We use the DeepScaleR dataset (Luo et al., 2025) as the training set. DeepScaleR compiles problems from multiple sources, including the American Invitational Mathematics Examination (AIME) from 1984 to 2023 and the American Mathematics Competitions (AMC) prior to 2023. The dataset also includes problems from the Omni-MATH (Gao et al., 2024) and Still datasets (Team, 2025), which feature problems from various national and international math competitions. This results in a diverse and challenging training set, covering a wide range of mathematical domains and difficulty levels.
@@ 4.3 Training Setup
We trained two models on the three difficulty-based distributions of the DeepScaleR dataset described in Section 4.2: Qwen 2.5 7B333https://huggingface.co/Qwen/Qwen2.5-7B and Qwen 2.5 MATH 1.5B444https://huggingface.co/Qwen/Qwen2.5-Math-1.5B. This setup allows us to evaluate the effectiveness of AdaRFT on models with different initial performance levels when exposed to skew-hard, skew-easy, and uniform problem distributions.
All models were trained using three different approaches: (1) the standard PPO algorithm, (2) AdaRFT (PPO), our method that integrates adaptive curriculum learning with PPO (see Section 3), and (3) PPO with filtered data, a strong baseline that trains PPO using filtered data based on pass@ accuracy. Following the approach of prior work (Bae et al., 2025; Hu et al., 2025; Zyphra, 2025), we perform a pass@40 analysis for each model and data distribution combination. For the data filtering baseline (3), we remove examples that are either too easy or too hard, excluding problems with solved percentages or . This focuses training on problems of intermediate difficulty. However, this filtering discards a significant portion of the training data, including many potentially useful examples. Moreover, because difficulty is determined via pass@ metrics, filtering must be recomputed each time the model or data distribution changes. In contrast, AdaRFT only requires difficulty estimation once per dataset and adapts to any model during training. For AdaRFT, the target difficulty was dynamically adjusted throughout training based on the model’s reward signal. This adjustment ensured that the model consistently encountered problems that matched its current skill level, preventing the model from being overwhelmed by overly difficult problems or stagnating on problems that were too easy.